OpenStudy (mrmanman305):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals four divided by x. and g of x equals four divided by x heeeeelp

Vocaloid (vocaloid):

well right now we have f(x) = 4/x and g(x) = 4/x find f(g(x))

OpenStudy (mrmanman305):

would you do 4(4/x)

Vocaloid (vocaloid):

should be 4/(4/x)

OpenStudy (mrmanman305):

sorry that sent twice

Vocaloid (vocaloid):

since f(x) = 4/x and g(x) = 4/x, replace "x" with "4/x" to get 4/(4/x)

OpenStudy (mrmanman305):

is that equal to x?

Vocaloid (vocaloid):

well, you tell me. do you remember the rule for dividing by fractions?

OpenStudy (mrmanman305):

oh yes. it would be equal to x lol

Vocaloid (vocaloid):

yup! f(g(x)) = x, and you can probably figure out that g(f(x)) also equals x since it's the same math process

OpenStudy (mrmanman305):

much appreciation @Vocaloid thank you